OpenStudy (noodya123):

CAN SOMEONE PLEASE HELP A ball is thrown from a height of 65 meters with an initial downward velocity of 10 m/s. The ball's height h (in meters) after t seconds is given by the following. h=65-10t-5t^2 How long after the ball is thrown does it hit the ground? Round your answer(s) to the nearest hundredth. (If there is more than one answer, use the "or" button.)

OpenStudy (kinged):

When the ball hits the ground h = 0 0 = 65-10t-5t^2 Factor out -5 0 = 5 ( t^2 + 2t - 13 ) Using the quadratic formula t ∈ { 2.7416573867739413 }

OpenStudy (kinged):

\[\approx2.74s\]

OpenStudy (noodya123):

i did the quaderatic formula but got 4 lol so i obviously did something wrong but thank you

OpenStudy (kinged):

You mean -4 which is cancelled.

OpenStudy (noodya123):

ya

OpenStudy (kinged):

You just take the other root which is positive since there's no such thing as a negative time (at least in this case).

OpenStudy (noodya123):

Can you help me with this one A ball is thrown from an initial height of 2meters with an initial upward velocity of 15m/s. The ball's height h (in meters) after t seconds is given by the following. h=2+15t-5t^2 Find all values of t for which the ball's height is 8 meters.